The variables x and y are related by the following equation: x = 3y2 - 5
If x = 103, what positive value of y makes this an equality?

Answer by richwmiller(17219)

x=3y^2-5
substitute 103 for x
103=3y^2-5
108=3y^2
12*9=3y^2
12*3=y^2
36=y^2
6=y